Minutes of nanoBPM meeting on 30 September, 2004

"Installation of the KEK system", H. Yamaoka

(transparencies, 5 pages, pdf, 305KB )

We installed the KEK NanoBPM system of about 2ton weight at the ATF extraction line for 9/27,9am through 9/28,9pm. The granite support table was fixed with anchor bolts on the floor. Alignment was completed by SMART (3D laser alignment system). The errors were estimated to be x <0.1mm, y<0.1mm, where z is beam axis and x, y are horizontal and vertical coordinate, respectively. We found that the beam level was 1205mm and 361mm from the floor and the granite table which was planned to be 1198mm and 371mm, respectively. So, the level was adjusted by mechanical jacks. Longitudinal position (z) of the granite table was off to the downstream by 13mm. Adjusting the z position of the system by the SUS plate on the granite table, finally the central cavity was set at 3mm upstream.

There are some problems as follows. Because of the z adjustment, the SUS plate could not bolted on the granite table. The SUS plate will be fixed by special metal studs. Extensions for the laser frame could not be attached to the reference bar for space limitation. Since the beam line is only 20mm off from the reference-bar support leg in x-direction, the diameter of bellows must be carefully checked.

"Status of KEK nano-BPM", Y. Inoue

First, resonance frequencies of all three cavity-BPMs were measured together with three reference cavities. The results (in unit of GHz) are listed in a following table.

nano-BPMno.1no.2no.3
reference cavity6.54046.53326.5400
sensor cavity 6.55116.55056.5515
Since the two frequencies must be coincident within 1MHz, the reference cavities will be adjusted by pushing three holes in each cavity-wall. The squeeze of cavities could increase the frequencies, while the frequencies could decrease by heating the cavities up. Coupling between two mutually-perpendicular output-ports was also measured to -20dB.

Previous measurement showed that the electrical center in the cavity was off 1mm from the mechanical center. As planned, the method with bead of larger diameter, i.e. 3mm&phi, was applied to the first cavity. Two resonant peaks were clearly observed with the bead set off from the electrical center, while the single peak was expected with the bead set at the center. Within 1mm around the center, only the single peak was observed. So, this method could not have necessary sensitivity to determine the electrical center. The antenna method will be applied to the second and third nano-BPM cavities. Also, the wire method will be tried, where transmitted electromagnetic waves are detected instead of reflected wave.

"Status of mover system", Y.Honda

(transparencies, xx pages, pdf, 1.4MB )

Three sets of V-blocks have been machined for bases of the nano-BPMs with precision of 10μm. During the machining, a voltage of 70V was applied to the Piezo devices, where the maximum voltage is 150V, and also vibrational amplitude was measured to be 12 μm. Alignment of the three BPMs was checked to be 10 μm three times by a dial gauge.

48 cables will be installed in the ATF-extraction from the roof concrete shield. A rack will be installed for the electronics.

The schedule was reminded as follows;

We agreed to have weekly meeting until the beam comes back at least. So, next meeting will be held in 7 October, 2004, at 1:30pm-, 3-gokan 325.